# Tackling Graph Oversquashing by Global and Local Non-Dissipativity

### v1

@@ Appendix I The Stability of the Jacobian
As discussed in Section 2.3, assuming that the Jacobian of the underlying system does not change significantly over time allows us to analyze the system from an autonomous system perspective (Ascher & Petzold, 1998) and mirrors prior approaches (Ruthotto & Haber, 2020; Chen et al., 2018a; Ruthotto & Haber, 2019; Chang et al., 2019; Chamberlain et al., 2021b; Gravina et al., 2023). In addition to building on existing literature, below, we provide an empirical measurement on a real-world dataset (pepties-func) of the Jacobian of our SWAN over time (layers). For reference, we compare it with the Jacobian of GCN. As can be seen from Figure 7, the Jacobian of SWAN has a minimal Jacobian change over time with an average of 0.6% between layers, while the change in the Jacobian over time in GCN is 40% on average.

### v2

@@ Appendix I The Stability of the Jacobian
As discussed in Section 2.3, assuming that the Jacobian of the underlying system does not change significantly over time allows us to analyze the system from an autonomous system perspective (Ascher and Petzold 1998) and mirrors prior approaches (Ruthotto and Haber 2020; Chen et al. 2018; Chang et al. 2019; Chamberlain et al. 2021b; Gravina, Bacciu, and Gallicchio 2023). In addition to building on existing literature, below, we provide an empirical measurement on a real-world dataset (peptides-func) of the Jacobian of our SWAN over time (layers). For reference, we compare it with the Jacobian of GCN. As can be seen from Figure 8, the Jacobian of SWAN has a minimal Jacobian change over time with an average of 0.6% between layers, while the change in the Jacobian over time in GCN is 40% on average.
@@
Appendix J Limitations and Broader Impact
This paper aims to contribute to the field of Machine Learning, specifically focusing on advancing Differential-Equation inspired Graph Neural Networks (DE-GNNs). The research presented herein has a positive impact on the ongoing exploration and applications of GNNs, with an emphasis on applications that require long-range propagation or suffer from oversquashing. Potentially, it might be used in other aspects of GNNs.
Although we believe that our work has a positive impact, it is important to note that SWAN, as our title suggests, aims to propose a mechanism to address the oversquashing problem in GNNs through the perspective of DE-GNNs, and it is therefore focused on problems that require long interaction modeling. Also, SWAN provides an understanding of the model and its expected behavior, while keeping low computational costs (as shown and analyzed in Appendix F), and also offers significant improvement in its class, as discussed in the paper. However, if one aims to achieve the highest downstream performance, regardless of the computational complexity and the theoretical understanding, other solutions, such as higher-order GNNs or transformers, may also be viable and achieve higher performance at times.